Circular No. 6609

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997bm IN IC 705 J. Mueller reports her discovery of a supernova (mag ∼ 18) on an IV

N plate taken on Mar. 18 by K. M. Rykoski and J. D. Mendenhall with the 1.2-m Oschin Schmidt Telescope in the course of the second Palomar Sky Survey. SN 1997bm is located 8′′ east and 13′′.4 north of the center of IC 705 (α = 11^{h}32^{m}.9, δ = +50^{o}14′, equinox 2000.0). A spectrogram obtained by W. Sargent and A. Baker on Mar. 30 at the 5-m Hale telescope (+ double spectrograph) indicates this to be a probable type-I supernova.

SUPERNOVA 1997bl IN ANONYMOUS GALAXY Further to the report on IAUC 6605, D. Hardin reports that the su-

pernova is located 2′′ northwest of the host galaxy’s center. EROS is an acronym for “Experience de Recherche d’Objets Sombres”, which searches for microlensing events and for supernovae.

COMET C/1995 O1 (HALE-BOPP) H. Matthews, Joint Astronomy Centre (JAC), Hilo, and National Re-

search Council of Canada; W. Holland, JAC; and R. Meier and D. Jewitt, University of Hawaii, report the first images of C/1995 O1 obtained with the SCUBA bolometer array at the James Clerk Maxwell Telescope on Mauna Kea: “Images taken at 850 μm on Mar. 30.92 UT show an approximately spherical coma that is extended relative to the 14′′ diffraction-limited resolution. Preliminary flux densities within synthetic apertures of projected diameters 14′′ and 60′′ are 720 ± 50 and 2180 ± 200 mJy, respectively, corresponding to equivalent blackbody cross-sections 2.5 ± 0.2 and 7.7 ± 0.7 × 10^{4} K km^{2} at assumed temperature 300 K. The surface-brightness profile appears slightly steeper than expected on the basis of steady-state ejection. The inferred central excess (210 mJy) places a strong upper limit of 49 km to the radius of the nucleus.”

Naked-eye m_{1} estimates: Mar. 19.46 UT, −0.6 (G. W. Kronk, Troy, IL); 22.14, −0.7 (R. J. Bouma, Groningen, The Netherlands); 25.03, −0.3 (J. Bortle, Stormville, NY, strong moonlight); 26.81, −1.2 (K. Černis, Zirmunai, near Vilnius, Lithuania); 27.85, −1.5 (A. Pereira, Cabo da Roca, Portugal); 29.87, −0.5 (H. Luethen, Telmer, Germany); 30.84, −0.3 (J. D. Shanklin, Cambridge, England); 31.85, −1.0 (J. Carvajal, Guadalajara, Spain); Apr. 1.80, −0.8 (M. V. Zanotta, Milan, Italy).

1997 April 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green